SpaceRef

SpaceRef


SOHO Daily Report Operations Day - 142/2003 (22 May 03)

Status Report From: Goddard Space Flight Center
Posted: Sunday, May 25, 2003


Mission Day 2729
SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC
Ref. #:  SH-FOT-DR-030522
====================================================================
PASS #1	Madrid			DSS66		0507-1540
PASS #2	Goldstone		DSS16		1728-2325
*********************************************************************

PASS #1	Madrid			DSS66		0507-1540

None

NRT:

1254-1300
1306-1311

Activities Summary:

- Collected MDI-M (0731-0831,1301-1521)
- Collected MDI-H (0513-0731, 0831-1301)
- Reaction Wheel Speed Monitoring Daily Adjustment
- CRS Drift Estimation Update
- SSR SEF Reset (x2)
- SSR Record
- New Tracking Star Slot 3 / Magnitude 7.15
- Hx Limit Update
- CAE-A in the AOCS HW Config Word

------------------

PASS #2	Goldstone		DSS16		1728-2325

Anomalies:

G03-0362  1735  Command failure due to TCP reboot (Telemetry & 
command outage for 2 minutes at the beginning of pass - data 
recoverable from SSR)  DR#G102783
G03-0363  2230  Degraded telemetry due to RFI (55 minutes of degraded 
data - 13 minutes is recoverable from SSR) DR#G102784

NRT:

2142-2237

Activities Summary:

- SSR Dump
- Collected MDI-M (1842-2305)
- Uplinked delayed load CEL030522301, SWA030522131
- Collected Ranging Data
- SSR SEF Reset (x1)
- SSR Record
------------------

-- End of Report --
-- 

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter